# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### v6

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets, and more than 10 different GLLMs together with 3 different prompting methods using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In 6 out of 7 cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $10$ percentage points.
@@ 4.2 Experimental Evaluation
Dataset Method gpt-4-o gpt-4-vision gemini-1.5-pro mistral-large claude-35
DocVQA Simple 0.656 0.759 0.647 0.445 0.653
Latin Prompting 0.718 - 0.668 0.447 0.739
SFT (Ours) 0.747 - 0.737 0.648 0.769
MPDocVQA Simple 0.686 0.708 0.627 0.364 0.581
Latin Prompting 0.718 - 0.726 0.335 0.561
SFT (Ours) 0.762 - 0.794 0.476 0.582
Kleister Charity Simple 0.711 0.751 0.764 0.652 0.774
Latin Prompting 0.721 - 0.763 0.576 0.786
SFT (Ours) 0.737 - 0.779 0.657 0.773
Kleister NDA Simple 0.649 0.664 0.722 0.637 0.623
Latin Prompting 0.646 - 0.707 0.624 0.626
SFT (Ours) 0.655 - 0.715 0.641 0.635
SROIE Simple 0.887 0.834 0.594 0.855 0.922
Latin Prompting 0.863 - 0.623 0.863 0.89
SFT (Ours) 0.897 - 0.623 0.905 0.917
VRDU AD Buy Simple 0.596 0.640 0.645 0.386 0.569
Latin Prompting 0.655 - 0.686 0.435 0.609
SFT (Ours) 0.753 - 0.765 0.594 0.612
VRDU Registration Simple 0.605 0.665 0.730 0.579 0.663
Latin Prompting 0.611 - 0.749 0.587 0.664
SFT (Ours) 0.605 - 0.780 0.639 0.67
For the evaluation, we always report the values of the newest model version. A detailed comparison between different intermediate versions of GPT as well as Gemini is given in the appendix. The following models were evaluated in this table:
  1. 1.
    gpt-3.5-turbo-16k (gpt-3.5-turbo-16k-0613)
  2. 2.
    gpt-4-turbo (Vgpt-4-1106-preview)
  3. 3.
    gpt-4-vision (gpt-4-vision-preview-1106) [2]
  4. 4.
    gemini-pro (gemini-1.5-pro-preview-0409) [17]
  5. 5.
    mistral-large (Feb. 2024) [18]
  6. 6.
    claude-3 (claude-3-opus) [16]
@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k-0613 gpt-3.5-turbo-0125 gpt-4-1106-preview gpt-4-turbo-2024-04-09 gpt-4-vision-preview-1106 gpt-4o-2024-05-13
DocVQA Simple 0.576 0.550 0.607 0.589 0.759 0.656
Latin Prompting 0.620 0.626 0.699 0.631 - 0.718
SFT (Ours) 0.768 0.733 0.790 0.770 - 0.747
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708 0.686
Latin Prompting 0.489 0.507 0.739 0.656 - 0.718
SFT (Ours) 0.724 0.731 0.781 0.760 - 0.762
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751 0.711
Latin Prompting 0.477 0.493 0.735 0.728 - 0.721
SFT (Ours) 0.534 0.548 0.763 0.733 - 0.737
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664 0.649
Latin Prompting 0.412 0.593 0.705 0.643 - 0.646
SFT (Ours) 0.315 0.562 0.703 0.659 - 0.655
SROIE Simple 0.509 0.706 0.835 0.849 0.834 0.887
Latin Prompting 0.723 0.708 0.851 0.847 - 0.863
SFT (Ours) 0.792 0.649 0.873 0.890 - 0.897
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640 0.596
Latin Prompting 0.424 0.211 0.586 0.506 - 0.655
SFT (Ours) 0.676 0.452 0.770 0.712 - 0.753
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665 0.605
Latin Prompting 0.606 0.630 0.673 0.614 - 0.611
SFT (Ours) 0.637 0.654 0.711 0.590 - 0.605
@@ Appendix B Results Gemini
Dataset Method gemini-1.0-pro-001 gemini-1.5-pro-preview-0409
DocVQA Simple 0.586 0.647
Latin Prompting 0.676 0.668
SFT (Ours) 0.741 0.737
MPDocVQA Simple 0.603 0.627
Latin Prompting 0.502 0.726
SFT (Ours) 0.616 0.794
Kleister Charity Simple 0.583 0.764
Latin Prompting 0.478 0.763
SFT (Ours) 0.633 0.779
Kleister NDA Simple 0.623 0.722
Latin Prompting 0.599 0.707
SFT (Ours) 0.552 0.715
SROIE Simple 0.263 0.594
Latin Prompting 0.371 0.623
SFT (Ours) 0.288 0.623
VRDU AD Buy Simple 0.510 0.645
Latin Prompting 0.556 0.686
SFT (Ours) 0.685 0.765
VRDU Registration Simple 0.699 0.730
Latin Prompting 0.740 0.749
SFT (Ours) 0.720 0.780
@@ Appendix C Results Claude
Dataset Method claude-3-opus-20240229 claude-3-5-sonnet-20240620
DocVQA Simple 0.768 0.653
Latin Prompting 0.762 0.739
SFT (Ours) 0.831 0.769
MPDocVQA Simple 0.636 0.581
Latin Prompting 0.438 0.561
SFT (Ours) 0.575 0.582
Kleister Charity Simple 0.800 0.774
Latin Prompting 0.787 0.786
SFT (Ours) 0.786 0.773
Kleister NDA Simple 0.673 0.623
Latin Prompting 0.670 0.626
SFT (Ours) 0.677 0.635
SROIE Simple 0.933 0.922
Latin Prompting 0.926 0.89
SFT (Ours) 0.949 0.917
VRDU AD Buy Simple 0.577 0.569
Latin Prompting 0.608 0.609
SFT (Ours) 0.633 0.612
VRDU Registration Simple 0.685 0.663
Latin Prompting 0.715 0.664
SFT (Ours) 0.705 0.67

### v9

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets, and more than 20 different GLLMs together with 3 different prompting methods using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In almost all cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $10$ percentage points.
@@ 4.2 Experimental Evaluation
Dataset Method gpt-4-o gpt-4.5 gemini-1.5-pro mistral-large claude-37 Llama-3.1-405B
DocVQA Simple 0.72 0.760 0.655 0.445 0.763 0.653
Latin Prompting 0.742 0.794 0.734 0.447 0.823 0.595
SFT (Ours) 0.809 0.855 0.752 0.648 0.834 0.737
MPDocVQA Simple 0.623 0.677 0.643 0.364 0.713 0.625
Latin Prompting 0.736 0.767 0.755 0.335 0.780 0.640
SFT (Ours) 0.735 0.777 0.794 0.476 0.790 0.753
Kleister Charity Simple 0.723 0.765 0.748 0.652 0.778 0.665
Latin Prompting 0.736 0.775 0.747 0.576 0.778 0.643
SFT (Ours) 0.752 0.775 0.764 0.657 0.786 0.680
Kleister NDA Simple 0.626 0.638 0.613 0.637 0.629 0.608
Latin Prompting 0.635 0.637 0.608 0.624 0.616 0.591
SFT (Ours) 0.625 0.641 0.623 0.641 0.64 0.610
SROIE Simple 0.905 0.855 0.686 0.855 0.935 0.853
Latin Prompting 0.912 0.836 0.672 0.863 0.877 0.779
SFT (Ours) 0.897 0.877 0.681 0.905 0.899 0.879
VRDU AD Buy Simple 0.645 0.668 0.545 0.386 0.625 0.453
Latin Prompting 0.721 0.725 0.579 0.435 0.639 0.546
SFT (Ours) 0.768 0.726 0.734 0.594 0.641 0.700
VRDU Registration Simple 0.628 0.664 0.549 0.579 0.658 0.614
Latin Prompting 0.628 0.667 0.555 0.587 0.669 0.632
SFT (Ours) 0.624 0.666 0.585 0.639 0.673 0.614
For the evaluation, we always report the values of the newest model version. A detailed comparison between different intermediate versions of GPT as well as Gemini is given in the appendix.
@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k gpt-3.5-turbo gpt-4 gpt-4-turbo gpt-4-vision gpt-4o gpt-4o + vision gpt-4o-mini gpt-4o gpt-4.5
Version 0613 0125 1106-preview 2024-04-09 preview-1106 2024-05-13 2024-05-13 2024-07-18 2024-08-06 preview-2025-02-27
DocVQA Simple 0.576 0.550 0.607 0.589 0.759 0.656 - 0.568 0.72 0.760
Latin Prompting 0.620 0.626 0.699 0.631 - 0.718 - 0.623 0.742 0.794
SFT (Ours) 0.768 0.733 0.790 0.770 - 0.747 0.754 0.754 0.809 0.855
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708 0.686 - 0.577 0.623 0.677
Latin Prompting 0.489 0.507 0.739 0.656 - 0.718 - 0.584 0.736 0.767
SFT (Ours) 0.724 0.731 0.781 0.760 - 0.762 0.789 0.731 0.735 0.777
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751 0.711 - 0.647 0.723 0.765
Latin Prompting 0.477 0.493 0.735 0.728 - 0.721 - 0.672 0.736 0.775
SFT (Ours) 0.534 0.548 0.763 0.733 - 0.737 0.749 0.707 0.752 0.775
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664 0.649 - 0.622 0.626 0.638
Latin Prompting 0.412 0.593 0.705 0.643 - 0.646 - 0.633 0.635 0.637
SFT (Ours) 0.315 0.562 0.703 0.659 - 0.655 0.648 0.628 0.625 0.641
SROIE Simple 0.509 0.706 0.835 0.849 0.834 0.887 - 0.863 0.905 0.855
Latin Prompting 0.723 0.708 0.851 0.847 - 0.863 - 0.836 0.912 0.836
SFT (Ours) 0.792 0.649 0.873 0.890 - 0.897 0.892 0.872 0.897 0.877
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640 0.596 - 0.398 0.645 0.668
Latin Prompting 0.424 0.211 0.586 0.506 - 0.655 - 0.436 0.721 0.725
SFT (Ours) 0.676 0.452 0.770 0.712 - 0.753 0.755 0.693 0.768 0.726
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665 0.605 - 0.594 0.628 0.664
Latin Prompting 0.606 0.630 0.673 0.614 - 0.611 - 0.608 0.628 0.667
SFT (Ours) 0.637 0.654 0.711 0.590 - 0.605 0.615 0.596 0.624 0.666
@@ Appendix B Results Gemini
Dataset Method gemini-1.0-pro gemini-1.5-pro-preview gemini-1.5-pro gemini-1.5-pro + vision gemini-1.5-flash gemini-2.0-flash
Version 001 0409 001 001 001 001
DocVQA Simple 0.586 0.647 0.655 - 0.64 0.689
Latin Prompting 0.676 0.668 0.734 - 0.707 0.773
SFT (Ours) 0.741 0.737 0.752 0.850 0.821 0.770
MPDocVQA Simple 0.603 0.627 0.643 - 0.696 0.680
Latin Prompting 0.502 0.726 0.755 - 0.665 0.680
SFT (Ours) 0.616 0.794 0.794 0.870 0.786 0.816
Kleister Charity Simple 0.583 0.764 0.748 - 0.720 0.735
Latin Prompting 0.478 0.763 0.747 - 0.732 0.753
SFT (Ours) 0.633 0.779 0.764 0.695 0.752 0.754
Kleister NDA Simple 0.623 0.722 0.613 - 0.592 0.622
Latin Prompting 0.599 0.707 0.608 - 0.596 0.623
SFT (Ours) 0.552 0.715 0.623 0.520 0.600 0.620
SROIE Simple 0.263 0.594 0.686 - 0.689 0.737
Latin Prompting 0.371 0.623 0.672 - 0.664 0.658
SFT (Ours) 0.288 0.623 0.681 0.931 0.699 0.683
VRDU AD Buy Simple 0.510 0.645 0.545 - 0.597 0.674
Latin Prompting 0.556 0.686 0.579 - 0.628 0.706
SFT (Ours) 0.685 0.765 0.734 0.545 0.701 0.746
VRDU Registration Simple 0.699 0.730 0.549 - 0.593 0.650
Latin Prompting 0.740 0.749 0.555 - 0.633 0.620
SFT (Ours) 0.720 0.780 0.585 0.570 0.667 0.613
@@ Appendix C Results Claude
Dataset Method claude-3-opus claude-3-5-sonnet claude-3-7-sonnet
Version 20240229 20240620 20250219
DocVQA Simple 0.768 0.653 0.763
Latin Prompting 0.762 0.739 0.823
SFT (Ours) 0.831 0.769 0.834
MPDocVQA Simple 0.636 0.581 0.713
Latin Prompting 0.438 0.561 0.780
SFT (Ours) 0.575 0.582 0.790
Kleister Charity Simple 0.800 0.774 0.778
Latin Prompting 0.787 0.786 0.778
SFT (Ours) 0.786 0.773 0.786
Kleister NDA Simple 0.673 0.623 0.629
Latin Prompting 0.670 0.626 0.616
SFT (Ours) 0.677 0.635 0.640
SROIE Simple 0.933 0.922 0.935
Latin Prompting 0.926 0.89 0.877
SFT (Ours) 0.949 0.917 0.899
VRDU AD Buy Simple 0.577 0.569 0.625
Latin Prompting 0.608 0.609 0.639
SFT (Ours) 0.633 0.612 0.641
VRDU Registration Simple 0.685 0.663 0.658
Latin Prompting 0.715 0.664 0.669
SFT (Ours) 0.705 0.670 0.673